A monopolist’s demand function is given as Q,=2000-10P where Q is the quantity is produced and sold and P is the price per unit in Ksh. If the firm’s marginal cost is K.sh100:

  1. Calculate the monopolist’s equilibrium quantity and price.   
1
Expert's answer
2021-05-09T14:49:19-0400

Monopolist demand function is

Q=2000 -10P

Inverse demand function is

"10P= 2000 -Q \\

P= 200 - \frac{Q}{10}"

Total revenue= Price "\times" Quantity

"= (200 - \frac{Q}{10})Q \\

= 200Q -\frac{Q^2}{10}"

"Marginal revenue = \frac{\partial TR}{\partial Q} \\

MR= 200 - 2 \times \frac{Q}{10} \\

= 200 - \frac{Q}{5}"

MC= 100 given

Monopolist equilibrium condition

"MR= MC \\

200 - \frac{Q}{5}=100 \\

\frac{Q}{5} = 200 -100 \\

\frac{Q}{5}= 100 \\

Q = 100 \times 5 \\

Q = 500"

Price equation

"P= 200 - \frac{Q}{10}"

Substituting Q into price equation

"P= 200 - \frac{500}{10} \\

P= 200 -50 \\

P= 150"

Equilibrium price will be 150 and quantity will be 500 units.